OpenStudy (anonymous):

The distance from your home to the local library is 5 1/4 miles. You are making a scale drawing of your neighborhood. How many inches need to be draw to represent the distance to the library using the scale of 1 inch=3 miles?

OpenStudy (anonymous):

A scale is a ratio, a ratio is a fraction and ratios are maintained so... 1inch=3miles But this is a ratio, that means that. inch/miles will be maintained. \[\frac{inches}{miles}=\frac{1}{3}=\frac{x}{5.25}\] So you can solve this: \[\frac{1}{3}=\frac{x}{5.25}\] And you've got your answer.

OpenStudy (anonymous):

thank you